OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 15x-5y=30

OpenStudy (saifoo.khan):

When you have to find the x-intercept, make y = 0 When you have to find the y-intercept, make x = 0

OpenStudy (anonymous):

15x-5y = 30 -5y = 30-15x y = 3x-6 y = 3(0)-6 y = 0-6 y = -6, (0,6) (0) = 3x-6 0 = 3x-6 -3x = -6 x = 2, (2,0) Solution: x-intercept = (2,0); y-intercept = (0,-6)

OpenStudy (anonymous):

for x intercept take y=0 then we will get x=2; for y intercept x=0 then we will get y=-6